# MinD-3D: Reconstruct High-quality 3D objects in Human Brain

### v1

@@ Abstract
In this paper, we introduce Recon3DMind, a groundbreaking task focused on reconstructing 3D visuals from Functional Magnetic Resonance Imaging (fMRI) signals. This represents a major step forward in cognitive neuroscience and computer vision. To support this task, we present the fMRI-Shape dataset, utilizing 360-degree view videos of 3D objects for comprehensive fMRI signal capture. Containing 55 categories of common objects from daily life, this dataset will bolster future research endeavors. We also propose MinD-3D, a novel and effective three-stage framework that decodes and reconstructs the brain’s 3D visual information from fMRI signals. This method starts by extracting and aggregating features from fMRI frames using a neuro-fusion encoder, then employs a feature bridge diffusion model to generate corresponding visual features, and ultimately recovers the 3D object through a generative transformer decoder. Our experiments demonstrate that this method effectively extracts features that are valid and highly correlated with visual regions of interest (ROIs) in fMRI signals. Notably, it not only reconstructs 3D objects with high semantic relevance and spatial similarity but also significantly deepens our understanding of the human brain’s 3D visual processing capabilities. Project page at: https://jianxgao.github.io/MinD-3D.
Figure 1: We introduce a novel task termed as Recon3DMind. Left: we showcase our data acquisition and modeling principles. Right: our proposed MinD-3D decodes fMRI signals to recover semantically matched, structurally coherent 3D objects with remarkable fidelity.
@@ 1 Introduction
Functional Magnetic Resonance Imaging (fMRI), a kind of signal that can be obtained in a non-invasive way, could capture blood changes in the human brain induced by neuronal activity. Due to its relatively easy accessibility, fMRI has been commonly used to reflect visual activities. Some recent studies [4, 5, 31, 27, 26] have successfully reconstructed high-quality images from fMRI signals either by utilizing powerful generative models e.g., diffusion models [29] or using generative transformers [3]. These approaches focus on extracting semantic features from fMRI signals, often requiring only semantic features to generate relevant high-quality images.
@@
Some literature [19, 13, 15, 20] has proved that the brain’s visual processing for 3D perception is far more complex than for 2D perception. Consequently, relying solely on semantic features is inadequate for accurately modeling the brain’s capability for 3D spatial imaging. To describe 3D objects effectively, it is essential to consider their shape and structural attributes in addition to their semantic properties. For instance, two cars of varying lengths might look identical from the front but show significant differences when viewed from the side. This observation underscores the importance of capturing the full range of spatial and structural features for a true representation of 3D objects. Therefore, our task focuses on enhancing modeling of human 3D comprehension. It calls for a sophisticated fMRI feature extractor that can discern not just semantic elements but also the spatial structure and other 3D-specific features from fMRI signals, thus facilitating a more complete and accurate recovery of 3D visual information.
Addressing the unique challenges of our Recon3DMind task, we confront a significant hurdle: the absence of a dataset pairing fMRI data with 3D visuals. To fill this void, we introduce fMRI-Shape, a pioneering dataset specifically designed for this purpose. Drawing inspiration from ZeroNVS [30], which posits that 3D objects can be effectively represented through 360-degree view videos, we adopt this methodology for fMRI data collection. In our approach, we present subjects in ShapeNet [2] with 360-degree view videos of stationary 3D objects, where the camera circumnavigates the object, completing a full rotation. This setup enables the capture of fMRI signals as subjects observe the objects from every conceivable angle. Such comprehensive visual exposure ensures the collection of detailed and precise fMRI signals, effectively capturing the entire spectrum of the object’s spatial attributes. Through intricate preprocessing steps, these recordings are transformed into multi-frame fMRI signals. The nuances and intricacies of the fMRI-Shape dataset will be elaborated further in Sec. 3.
@@
The second stage focuses on decoding the brain’s visual activity features from the extracted fMRI data. We achieve this through a diffusion model trained within the feature space, conditioned on the fMRI features. This conditioning allows us to progressively generate accurate visual representations, effectively translating complex brain activity into comprehensible visual data.
@@
To evaluate the efficacy of our model, we use both semantic and structural metrics. These metrics provide a comprehensive evaluation of the model’s ability to create structurally accurate 3D representations. Additionally, we validate the relevance of our model’s extracted features by correlating them with specific brain regions, ensuring that our model not only generates accurate representations but also does so in a way that is consistent with how the brain processes visual information.
@@
  • We for the first time define the novel task, Recon3DMind, which is meaningful for both computer vision and cognitive neuroscience communities.
  • We introduce an innovative, versatile brain decoding framework, composed of three distinct modules: a neuro-fusion encoder, a feature bridging diffusion model, and a latent adapted decoder.
  • We present fMRI-Shape, the first dataset of fMRI-3D shape pairs, to foster further research in this domain. This dataset will be accessible on our project page.
@@ 2.1 FMRI Decoding Methods
Current fMRI decoding methods primarily aim to reconstruct the real world in a 2D format, such as the images or videos that humans perceive. This task is quite complex, as it involves accurately extracting relevant features from fMRI signals and reconstructing a precise 2D representation. Deep learning methods, known for their impressive capabilities, are well-suited for addressing this challenge. Initial successes in this area have been demonstrated by earlier methods [18, 37, 33]. Subsequent studies [32, 7] have shown that generative models are more effective for these tasks. As a result, various diffusion models [4, 5, 31] have been employed as decoders to reconstruct visual scenes, achieving remarkable results. However, these studies have been limited to 2D visual representations. In this paper, we aim to extend the scope of fMRI decoding to 3D representations. Our goal is to directly reconstruct 3D objects from fMRI signals. To accomplish this, we propose a new framework that uses a transformer-based feature encoder for extraction and aggregation. This framework translates neural space data into visual space and employs a powerful 3D decoder to reconstruct the 3D object, utilizing features from the visual space.
2.2 Diffusion Model
Diffusion models [17, 34] are exceptional generative tools for both pixel and feature generation. As a variant, the latent diffusion model [29], equipped with an autoencoder, compresses images into lower-dimensional latent features, thereby generating a compressed version of the data rather than directly generating the data. This approach, which operates in the latent data space, significantly reduces computational requirements and enables the generation of higher-quality images with enhanced details in the latent space. In this paper, we aim to leverage the potent feature generation capabilities of diffusion models to generate visual features based on fMRI features. To achieve this, we have adapted the latent diffusion model by discarding its autoencoder structure and focusing solely on its latent component. The conditional information driving the model is derived from the fMRI features.
@@ 2.2 Diffusion Model
Advancements in deep generative models [36, 41] and 3D-aware image synthesis techniques [6, 38] have significantly influenced the field of 3D object generation. These models are instrumental in creating complex 3D shapes and structures, especially in architecture and virtual environment design, emphasizing discrete voxel generation and 3D modeling from 2D images. In our study, we aim to generate 3D objects from fMRI data, leveraging Argus [40], a robust 3D generative model, as our decoder and adapt it with several transformer layers. This approach facilitates an integrated method that combines fMRI and visual features produced by the previous module. The visual features serve as conditional embeddings for Argus, while the fMRI features are fused in the adapter layers. This synergistic integration aims to enhance the model’s capability to accurately reconstruct 3D objects from complex brain activities.
Figure 2: Overview of the fMRI-Shape Acquisition Process. Initially, we render each object into an 8-second long video, showcasing a 360-degree view. Following this, we capture fMRI signals in video format. The fMRI data are then processed using fMRIPrep, converting the data from 32k fs LR surface space into 2D images with dimensions of 1023 × 2514.
@@ 3 Experimental Designs and Curated Dataset
In this section, we will introduce the detailed procedure for the collection of the proposed dataset, fMRI-Shape. This dataset will be released to the community to push the research on the topic of Recon3DMind.
@@ 3.1 Subjects and Experiments
Three participants (male, aged 24-27), who were unaware of our work, were instructed to view videos depicting the gradual 360-degree rotation of 3D objects at a 60-degree pitch angle, as shown in Fig. 2. All participants had normal or corrected-to-normal visual acuity. Informed written consent was obtained from each participant, and the experimental protocol was approved by the ethical review board. To ensure the diversity of the dataset for our work, the 3D objects used in this study were derived from the ShapeNetCore [2], which includes 55 classes of objects represented as 3D point clouds, 2D images, and CAD models. We randomly selected four objects from each class in ShapeNetCore, allocating three for training and one for testing. We utilized the rendering technique from Zero123 [21] to generate an 8-second long video for each object in Blender.
For each participant, the experiment was divided into six runs. During each run, participants consecutively viewed 55 videos in a randomized order, one from each category, resulting in an approximate run duration of 8 minutes. The runs alternated between training and test data, with a 1-minute rest period after each run. Notably, each participant underwent three viewings of the testing run and a single viewing of the three training runs.
@@ 3.2 Data Acquisition and Preprocessing
MRI data were collected at Zhangjiang International Brain Imaging Centre. T1, T2 and fMRI data were acquired in a 3T scanner and a 32-channel RF head coil. T1-weighted data were scanned using MPRAGE sequence (0.8-mm isotropic resolution, TR=2500ms, TE=2.22ms, flip angle $8^{\circ}$). T2-weighted data were acquired using SPACE sequence 1(0.8-mm isotropic resolution, TR=3200ms, TE=563ms, flip angle $120^{\circ}$). Functional data were scanned using gradient-echo EPI at 2-mm isotropic resolution with whole-brain coverage (TR=800ms, TE=37ms, flip angle $52^{\circ}$, multi-band acceleration factor 8). The sampling frequency of the 3T scanner is 8Hz, so each video corresponds to a total of 10 frames of task-state fMRI signals.
@@
Preprocessing was performed using fMRIPrep ([10, 9]). Following [27], the preprocessed functional data in 32k_fs_LR surface space were converted into 2D images and utilized for further analysis. Specifically, for each 32k_fs_LR data frame in each run, we applied z-scoring to the values across every vertex within the run. These z-scored values were then mapped onto 1023 × 2514 2D images using pycortex [11]. The ROIs for analysis were selected from the HCP-MMP atlas [12] in the 32k_fs_LR space, including “V1, V2, V3, V3A, V3B, V3CD, V4, LO1, LO2, LO3, PIT, V4t, V6, V6A, V7, V8, PH, FFC, IP0, MT, MST, FST, VVC, VMV1, VMV2, VMV3, PHA1, PHA2, PHA3”.
@@ 4 Problem Setup
Figure 3: Overview of the MinD-3D Framework. Our approach combines a Neuro-Fusion Encoder for extracting features from fMRI frames, a Feature Bridge Diffusion Model for generating visual features from these fMRI signals, and a Latent Adapted Decoder based on the Argus 3D shape generator for reconstructing 3D objects. This integrated system effectively aligns and translates brain signals into accurate 3D visual representations. Note that the CLIP encoder is only for training the model, while it is not used for the inference stage.
@@
Consider the acquisition of a multi-frame fMRI signal denoted as $\{F\}$, where $|F|=10$. These signals correspond to both a 3D object point cloud, $\{P\}$, and a video, $\{V\}$, with $|V|=192$ frames, which the subject views. The task necessitates an efficient encoder, $E$, capable of extracting both spatial structural and semantic features from the fMRI signal. It’s crucial to note that while a single frame of the fMRI signal suffices for semantic information extraction in 2D image reconstruction, the reconstruction of 3D structures demands additional spatial structural features. Therefore, we input multiple frames of fMRI signals with the aim of extracting these comprehensive spatial structural features from the spatio-temporal data. This process is mathematically represented as: $f=E(F)$ Following the feature extraction, we employ a powerful decoder to reconstruct the original 3D point cloud $\{P\}$ based on the extracted feature $f$: $P=D(f)$ Consequently, our model can be succinctly defined as $M=\{E,D\}$, thereby operationalizing the transformation: $P=M(F)$.

### v3

@@ Abstract
In this paper, we introduce Recon3DMind, an innovative task aimed at reconstructing 3D visuals from Functional Magnetic Resonance Imaging (fMRI) signals, marking a significant advancement in the fields of cognitive neuroscience and computer vision. To support this pioneering task, we present the fMRI-Shape dataset, which includes data from 14 participants and features 360-degree videos of 3D objects to enable comprehensive fMRI signal capture across various settings, thereby laying a foundation for future research. Furthermore, we propose MinD-3D, a novel and effective three-stage framework specifically designed to decode the brain’s 3D visual information from fMRI signals, demonstrating the feasibility of this challenging task. The framework begins by extracting and aggregating features from fMRI frames through a neuro-fusion encoder, subsequently employs a feature bridge diffusion model to generate visual features, and ultimately recovers the 3D object via a generative transformer decoder. We assess the performance of MinD-3D using a suite of semantic and structural metrics and analyze the correlation between the features extracted by our model and the visual regions of interest (ROIs) in fMRI signals. Our findings indicate that MinD-3D not only reconstructs 3D objects with high semantic relevance and spatial similarity but also significantly enhances our understanding of the human brain’s capabilities in processing 3D visual information. Project page at: https://jianxgao.github.io/MinD-3D.
Keywords:
fMRI 3D vision diffusion model
@@ 1 Introduction
Functional Magnetic Resonance Imaging (fMRI), a kind of signal that can be obtained in a non-invasive way, could capture blood changes in the human brain induced by neuronal activity. Due to its relatively easy accessibility, fMRI has been commonly used to reflect visual activities. Some recent studies [6, 7, 41, 36, 34] have successfully reconstructed high-quality images from fMRI signals by utilizing powerful generative models [39, 3]. These approaches focus on extracting semantic features from fMRI signals, often requiring only semantic features to generate relevant high-quality images.
Figure 1: Overview of Recon3DMind task, showcasing the fMRI-Shape dataset collection process with 14 participants observing 360-degree view videos of 3D objects, and MinD-3D framework for reconstructing 3D objects from fMRI signals.
@@
Some literature [15, 16, 25] has shown that the brain’s visual processing mechanisms for 3D perception are significantly more complex than those for 2D perception. This increased complexity is manifested in the differential activation of brain regions during 3D visualization tasks [14, 23]. Consequently, a reliance solely on semantic features falls short of accurately modeling the brain’s capacity for 3D spatial imaging. Effective description of 3D objects necessitates the consideration of their shape and structural attributes, in addition to semantic properties. For example, two cars may appear identical when viewed from the front, yet present substantial differences in length when observed from the side. This highlights the critical need to capture the full range of spatial and structural features for an authentic representation of 3D objects. Accordingly, our work aims to advance the modeling of human 3D comprehension. It necessitates the development of an advanced fMRI feature extractor capable of discerning not only semantic elements but also spatial structures and other 3D-specific characteristics from fMRI signals. This approach aims at enabling a more complete and accurate reconstruction of 3D visual information.
Addressing the unique challenges of our Recon3DMind task, we confront a significant hurdle: the absence of a dataset that pairs fMRI data with 3D visuals. To fill this void, we introduce fMRI-Shape, a pioneering resource specifically designed for this purpose, featuring data from 14 participants and supporting various experimental settings. Inspired by ZeroNVS [40], which suggests that 3D objects can be effectively represented through 360-degree view videos, we adopt a similar methodology for our fMRI data collection. In our approach, participants are presented with 360-degree view videos of stationary 3D objects from ShapeNet [2]. These videos feature the camera orbiting the object, completing a full rotation, thereby providing subjects with a comprehensive view of the objects from all possible angles. This method ensures the capture of detailed and precise fMRI signals as subjects observe the objects, effectively encapsulating the full range of the object’s spatial attributes. Through meticulous preprocessing, these recordings are converted into multi-frame fMRI signals, capturing a rich dataset that allows for a nuanced analysis. The complexities and specific details of the fMRI-Shape dataset will be further elaborated in Sec. 3.
@@
The second stage focuses on decoding the brain’s visual activity features from the extracted fMRI data. We achieve this through a transformer-based diffusion model trained within the feature space, conditioned on the fMRI features. This conditioning allows us to progressively generate accurate visual representations, effectively translating complex brain activity into comprehensible visual data.
@@
To evaluate the efficacy of our model, we employ various semantic and structural metrics. These metrics offer a comprehensive assessment of the model’s ability to generate structurally accurate 3D representations. Additionally, we validate the relevance of the features extracted by our model by correlating them with specific brain regions. This ensures that our model not only produces accurate representations but also does so in a manner consistent with the brain’s processing of visual information.
@@
  • For the first time, we propose and demonstrate the feasibility of the novel task, Recon3DMind, which holds significance for both the computer vision and cognitive neuroscience communities.
  • We present fMRI-Shape, the first dataset of fMRI-3D shape pairs, across various settings, to foster further research in this domain. This dataset will be accessible on our project page.
  • We introduce an innovative, versatile brain decoding framework MinD-3D, composed of three distinct modules: a neuro-fusion encoder, a feature-bridging diffusion model, and a latent-adapted decoder.
@@ 2.1 FMRI Decoding Methods
Current fMRI decoding methods primarily focus on reconstructing the vision perception in a 2D format, such as the images or videos perceived by humans. This task is complex, as it requires accurately extracting relevant features from fMRI signals and reconstructing a precise 2D representation. Deep learning methods, known for their impressive capabilities, are well-suited to address this challenge. Initial successes in this area have been demonstrated by earlier methods [19, 51, 43]. Subsequent studies [42, 10] have shown that generative models are particularly effective for these tasks, leading to the employment of various diffusion models [6, 7, 41, 45] as decoders to reconstruct visual scenes, achieving remarkable results. However, these studies have been limited to 2D visual representations and the related vision ROIs. In this paper, we aim to extend the scope of fMRI decoding to 3D representations, involving more vision ROIs. Our goal is to directly reconstruct 3D objects from fMRI signals. To accomplish this, we propose a new framework that employs a transformer-based feature encoder for extraction and aggregation. This framework translates neural space data into visual space and utilizes a powerful 3D decoder to reconstruct the 3D object, leveraging features from the visual space.
2.2 Diffusion Models
Diffusion models [18, 44] are exceptional generative tools for both pixel and feature generation. As a variant, the latent diffusion model [39], equipped with an autoencoder, compresses images into lower-dimensional latent features, thereby generating a compressed version of the data rather than directly generating the data itself. Dit [33] replaces the backbone of diffusion models with transformers, which will improve the performance and scalability of these models. This approach, operating in the latent space, significantly reduces computational requirements and enables the generation of higher-quality images with enhanced details in the latent space. In this paper, we aim to leverage the potent feature-generation capabilities of diffusion models to generate visual features based on fMRI features. To achieve this, we have adapted a transformer-based diffusion model, focusing solely on its latent component. The conditional information driving the model is derived from the fMRI features.
@@ 2.2 Diffusion Models
3D generation can be accomplished through various methods [35, 56, 48, 26]. Some methods [49, 48] employ 3D Gaussian splatting [24] for this purpose. Other studies [50, 54] utilize diffusion models to generate multi-view representations of objects, subsequently constructing 3D models. Additionally, traditional and direct approaches leverage autoregressive methods [8, 21, 35] for 3D object generation. In our study, we aim to generate 3D objects from fMRI data, using Argus [35], a robust 3D generative model, as our decoder, which we enhance with several transformer layers. This approach integrates fMRI and visual features produced by the preceding module. The visual features act as conditional embeddings for Argus, while the fMRI features are fused within the adapter layers. This synergistic integration seeks to augment the model’s capability to accurately reconstruct 3D objects from complex brain activities.
Figure 2: Comparing fMRI-Shape with other 2D fMRI datasets. As the first 3D fMRI dataset, fMRI-Shape features a larger number of participants and frames, providing ample support for experiments in our proposed novel task and further research.
Table 1: Details of fMRI-Shape Dataset. It ensures a balanced representation of male and female participants. Specifically, the Across-Person Set (AP Set) includes fMRI data sampled from different participants for Core test data. The Across-Person & Across-Class Set (APAC Set) involves different participants viewing new 3D objects. These two subsets facilitate evaluations for model generalization.
Participant Males/Females Category Objects Frames
Core Set 8 4/4 13 1404 14040
AP Set 2 1/1 13 104 1040
APAC Set 4 2/2 55 220 2200
Total 14 7/7 55 1624 123200
@@ 3 Experimental Designs and Curated Dataset
In this section, we will introduce the detailed procedure for collecting the proposed dataset, fMRI-Shape. The scale of fMRI-Shape compared to other datasets (NSD [1], BOLD5000 [4], GOD [19], Video-fMRI [51]) is illustrated in Fig. 2, while the specific details of fMRI-Shape are provided in Tab. 1. This dataset are released to the community to push the research on the topic of Recon3DMind.
Figure 3: Overview of the fMRI-Shape Acquisition Process. Initially, we render each object into an 8-second long video, showcasing a 360-degree view. Subsequent fMRI signal capture is performed in video format, followed by data processing with fMRIPrep to convert signals from 32k_fs_LR surface space into 2D images of dimensions 1023 $\times$ 2514. Individual differences observed in the dataset, as highlighted in the middle part, underscore the challenges in generalizing these findings. On the rights, regions of interest (ROIs) are transformed into 256 $\times$ 256 image.
@@ 3.1 Subjects and Experiments
FMRI-Shape encompasses data from 14 participants unfamiliar with the objectives of our study. All participants possessed normal or corrected-to-normal visual acuity. Informed written consent was obtained from each participant, and the experimental protocol received approval from the ethical review board.
To ensure the diversity of the dataset, the 3D objects used in this study were derived from the ShapeNetCore [2], which includes 55 classes of objects. We utilized the rendering technique from Zero123 [27] to generate an 8-second video for each object using Blender. These videos illustrate the gradual 360-degree rotation of 3D objects at a 60-degree pitch angle, as exemplified in Fig. 3.
Core The core set of fMRI-Shape includes data from 8 participants: 4 males and 4 females, aged 21 to 29 (Participants No. 1-8). We selected a total of 1404 objects across 13 common categories frequently employed in various 3D reconstruction literature [5, 46, 22] within ShapeNetCore. For each category, 100 objects were allocated for training and 8 for testing, totaling 108 objects per category. The experiment lasted approximately 4.5 hours per participant, divided into 32 sessions. In each session, participants viewed 48 videos in a randomized order, interspersed with 2.4 second rest periods, covering one object from each category, resulting in a session duration of around 9 minutes. Sessions alternated between training and test data, with a 1-hour break after every 8 sessions to prevent fatigue and maintain data quality. Notably, the testing objects were viewed twice by each participant.
Across-Person Set (AP Set) The AP Set, designed for Out-Of-Distribution (OOD) testing, includes fMRI data from 2 participants (1 male aged 24 and 1 female aged 26, Participants No. 9 and 10), sampled while they were viewing the test objects from the Core set.
Across-Person & Across-Class Set (APAC Set) The APAC Set represents a more challenging component than the AP Set. It includes data from 4 participants: 2 males and 2 females, aged 22 to 26 (Participants No. 11-14). We randomly selected 4 objects from each of the 55 categories in ShapeNetCore, distinct from those in the Core, for a total of 220 objects. Each participant was exposed to these 220 objects across 6 sessions.
As shown in the middle part of Fig. 3, we can observe individual differences, which will pose significant challenges for generalization. Thus, the AP and APAC sets are crucial for OOD testing and will serve as important benchmarks to assess the generalization capabilities of the 3D decoding model.
@@ 3.2 Data Acquisition and Preprocessing
T1 and fMRI data were acquired in a 3T scanner and a 32-channel RF head coil. T1-weighted data were scanned using MPRAGE sequence (0.8-mm isotropic resolution, TR=2500ms, TE=2.22ms, flip angle $8^{\circ}$). Functional data were scanned using gradient-echo EPI at 2-mm isotropic resolution with whole-brain coverage (TR=800ms, TE=37ms, flip angle $52^{\circ}$, multi-band acceleration factor 8). The sampling frequency of the 3T scanner is 1.25Hz, so each video segment corresponds to a total of 10 frames of task-state fMRI signals.
@@
Preprocessing was performed using fMRIPrep [12, 11]. Following [36], the preprocessed functional data in 32k_fs_LR surface space were converted into 2D images and utilized for further analysis. Given the delay of the BOLD signal by 6 seconds, we applied z-scoring to the data points across every vertex within each run, incorporating a 6-second lag. These normalized values were then projected onto 1023 × 2514 pixel 2D images using pycortex. For analysis, Regions of Interest (ROIs) were selected from the Human Connectome Project Multi-Modal Parcellation (HCP-MMP) atlas in the 32k_fs_LR space. These ROIs included areas such as “V1, V2, V3, V3A, V3B, V3CD, V4, LO1, LO2, LO3, PIT, V4t, V6, V6A, V7, V8, PH, FFC, IP0, MT, MST, FST, VVC, VMV1, VMV2, VMV3, PHA1, PHA2, PHA3”. Subsequently, the ROIs were converted into a 256 × 256 image, as illustrated in the right part of Fig. 3.
@@ 4 Problem Setup
Consider the acquisition of a multi-frame fMRI signal denoted as $\{F\}$, where $|F|=10$. These signals correspond to both a 3D object mesh, $\Psi$, and a video, $\{V\}$, with $|V|=192$ frames, which the subject views. The task necessitates an efficient encoder, $E$, capable of extracting both spatial structural and semantic features from the fMRI signal. It’s crucial to note that while a single frame of the fMRI signal suffices for semantic information extraction in 2D image reconstruction, the reconstruction of 3D structures demands additional spatial structural features. Therefore, we input multiple frames of fMRI signals with the aim of extracting these comprehensive spatial structural features from the spatio-temporal data. This process is mathematically represented as: $f=E(F)$ Following the feature extraction, we employ a powerful decoder to reconstruct the original 3D mesh $\Psi$ based on the extracted feature $f$: $\Psi=D(f)$ Consequently, our model can be succinctly defined as $M=\{E,D\}$, thereby operationalizing the transformation: $\Psi=M(F)$.